OpenStudy (anonymous):

What is the equation of a line passing through the point (a, b) and having a slope of b? y = −bx − b(1 − a) y = bx − b(a − 1)

OpenStudy (anonymous):

y=1/bx+b(a-1)

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

This is simple

OpenStudy (anonymous):

y=-1/bx+b(a+1)

OpenStudy (anonymous):

hahah other choices im so lost;-;

OpenStudy (anonymous):

I know

OpenStudy (anonymous):

Give me a second

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Wait, is this FLVS?

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

i used the slope point form and i actually got y-b=b(x-a) which if you get y by itself you would have y=b(x-a)+b

OpenStudy (anonymous):

Spot on

OpenStudy (anonymous):

thats not a choice ;-; i solved it and i dint get any of thoes

OpenStudy (anonymous):

y=bx-(ab)+b

OpenStudy (anonymous):

Those*

OpenStudy (anonymous):

then what is it, if you have already solved it?

OpenStudy (anonymous):

i got something wayyy diffrent lol i know its wrong

OpenStudy (anonymous):

Is there a four choice?

OpenStudy (anonymous):

im a science person not a math person

OpenStudy (anonymous):

Yes, but they are one in the same!

OpenStudy (anonymous):

so then what are alll the answers? list them all please. and what did you get?

OpenStudy (anonymous):

yeah i typed out the last two choices there in the comments

OpenStudy (anonymous):

hmm let me try something

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i tried it and i still have no idea. what class is this for?

OpenStudy (anonymous):

algerbra 1

OpenStudy (anonymous):

are there any numbers given to represent the letters?

OpenStudy (anonymous):

nope none

OpenStudy (anonymous):

Can I just give you the answer?

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

It was the one you listed in the comments

OpenStudy (anonymous):

i listed two

OpenStudy (anonymous):

First one

OpenStudy (anonymous):

wich one first one or last one?

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

Let me double check this

OpenStudy (anonymous):

its the second one of the first two you listed. i subsituted a set of numbers for a and b and the second one is the only one that crossed the point y = bx − b(a − 1)

OpenStudy (anonymous):

Yes, he is right. Sorry, I know, what I did wrong

OpenStudy (anonymous):

i used 2 for a and 4 for b and when you graph the equations that is the one that works

OpenStudy (anonymous):

its ok i have 94 questions left :(

OpenStudy (anonymous):

Oh my gosh

OpenStudy (anonymous):

Is this like a test or something

OpenStudy (anonymous):

no its a per test on plato

OpenStudy (anonymous):

Oh okay, I can see you are going to be there for a while

OpenStudy (anonymous):

yeah :(